OpenStudy (anonymous):

Help please!!! Simplify the following expression x-4y+5z-3x-y+2z a. -2x-5y-7z b.-2x-5y+7z c.-2z+5y-3z

OpenStudy (cggurumanjunath):

group all the x ,y and z terms .

OpenStudy (cggurumanjunath):

(x-3x) + (-4y-y) + (5z+2z)

OpenStudy (cggurumanjunath):

x(1-3)+y(-4-1)+z(5+2)

OpenStudy (cggurumanjunath):

is it clear ? @Coralgirl

OpenStudy (anonymous):

not quite but I am trying to understand it

OpenStudy (anonymous):

\(x-4y+5z-3x-y+2z=-2x-5y+7z\)

OpenStudy (anonymous):

\(x\)=apple apple-3apple=-2apple

OpenStudy (anonymous):

y=orange -4orange-orange=-5orange

OpenStudy (anonymous):

z=banana 5banana+2banana=7banana

OpenStudy (anonymous):

okay how to do you arrange it @AntarAzri ?

OpenStudy (anonymous):

how to do you rearranged them to solve ?

OpenStudy (anonymous):

-2apple-5orange+7banana and apple x, orange y, banana z

OpenStudy (anonymous):

What do you mean by arranging ?

OpenStudy (anonymous):

is it like -y-4y +5z+2z+3x-x

OpenStudy (anonymous):

Not like but it is

OpenStudy (anonymous):

btw, you must begin by apple, orange, then banana

OpenStudy (anonymous):

Okay ?

OpenStudy (anonymous):

-4y-y-3x-x+5z+2z?

OpenStudy (anonymous):

Well, you want one x one y and one z, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then, do the sum of your \(x\)s \(y\)s and \(z\)s

OpenStudy (anonymous):

x-3x = -2x

OpenStudy (anonymous):

Awsome

OpenStudy (anonymous):

What about y

OpenStudy (anonymous):

Come on do the sum

OpenStudy (anonymous):

-4y-y ?

OpenStudy (anonymous):

5y

OpenStudy (anonymous):

it's kinda \(y(-4-1)=?\)

OpenStudy (anonymous):

How much is \(-4-1=?\)

OpenStudy (anonymous):

last one is 7z

OpenStudy (anonymous):

Nope, you must get y first

OpenStudy (anonymous):

3

OpenStudy (anonymous):

sorry i mean \(4+1\) ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

that's right, \(-4y-y=-y(4+1)\) and that equal to what ?

OpenStudy (anonymous):

0kay

OpenStudy (anonymous):

Equal to what ?

OpenStudy (anonymous):

5y

OpenStudy (anonymous):

\((4+1)\) times \(-y\) ?

OpenStudy (anonymous):

\(4+1\) is \(5\) and \(-y\) that give us \(-5y\), right ?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

Did you studied the negative numbers ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am not getting this picture quite well. its kind of messy

OpenStudy (anonymous):

You come down \(4\) times from \(0\) which \(-4y\) and you come down again one more time which is \(-(1)y\) and you will be in the -5 degree, right

OpenStudy (anonymous):

Yup it's messy sorry

OpenStudy (anonymous):

yes, course

OpenStudy (anonymous):

no problem at all, I got it

OpenStudy (anonymous):



OpenStudy (anonymous):

Think of it that way

OpenStudy (anonymous):

Well, what if you come down 6 times and than 7 times where you will be ?

OpenStudy (anonymous):

-6, -7

OpenStudy (anonymous):

Yup the sum I mean ?

OpenStudy (anonymous):

Consider your self siting in \(0\) than you come down 6 times than 7 times again where you will be ?

OpenStudy (anonymous):

I mean you are siting on line :D

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

(SITTING)

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

-7

OpenStudy (anonymous):



OpenStudy (anonymous):

Right you are in the (-7)

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Sorry, I'm wrong, could you start from -6 and you come down 7 times ?

OpenStudy (anonymous):

Where you will be ?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

sorry -12

OpenStudy (anonymous):

Wrong, let me explain again

OpenStudy (anonymous):

sorry for trouble :(

OpenStudy (anonymous):

Are you sure -12 ?

OpenStudy (anonymous):

no i am not sure

OpenStudy (anonymous):

Isn't -13 ?

OpenStudy (anonymous):

Oh yea, i made mistake on counting